OpenStudy (anonymous):

The area of a triangle is 124 square units. What would its new area be if its base was half as long and its height was three times as long?

OpenStudy (anonymous):

I got 186, but I don't know if it's right.

OpenStudy (raden):

old area = 1/2 * b * h = 124 new area = 1/2 * (b/2) * (3h) = 3/2 * (1/2 * b * h) = 3/2 * 124 = ... ?